# Consistent-1-to-3: Consistent Image to 3D View Synthesis via Geometry-aware Diffusion Models

### v1

@@ Abstract
Zero-shot novel view synthesis (NVS) from a single image is an essential problem in 3D object understanding. While recent approaches that leverage pre-trained generative models can synthesize high-quality novel views from in-the-wild inputs, they still struggle to maintain 3D consistency across different views. In this paper, we present Consistent-1-to-3, which is a generative framework that significantly mitigate this issue. Specifically, we decompose the NVS task into two stages: (i) transforming observed regions to a novel view, and (ii) hallucinating unseen regions. We design a scene representation transformer and view-conditioned diffusion model for performing these two stages respectively. Inside the models, to enforce 3D consistency, we propose to employ epipolor-guided attention to incorporate geometry constraints, and multi-view attention to better aggregate multi-view information. Finally, we design a hierarchy generation paradigm to generate long sequences of consistent views, allowing a full 360 observation of the provided object image. Qualitative and quantitative evaluation over multiple datasets demonstrate the effectiveness of the proposed mechanisms against state-of-the-art approaches. Our project page is at https://jianglongye.com/consistent123/.
@@ 1 Introduction
1) Single-view $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
2) Sparse-views (2-4) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
3) Multi-view consistent $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
4) Generate unseen $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
5) Open-set generalization $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
6) Train-free for new instances $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 1: Comparison with prior methods. The rows indicate whether each method: 1) works with a single input view, 2) works with sparse (2-4) input views, 3) generate multi-view consistent views, 4) hallucinates unseen regions, 5) generalizes to open-set instances instead of limited categories, 6) free of time-consuming training for new instances. Our proposed method stands out as the only approach that possesses all these advantages.
@@
Synthesizing high-quality and visual consistent novel views of real-world objects from a single input image has been a long-standing problem in the computer vision community. It has a wide range of applications including content creation [17, 29], robotic manipulation and navigation [25], and AR/VR [14], etc.
@@
This task is challenging as it consists of difficulties form both reconstruction and generation, which requires not only accurate geometry transformation of observed parts, but also the hallucination of unseen regions. In recent years, the field has seen significant progress with the advance of large scale data-driven deep 2D models. For example, transformer models for single view reconstruction, e.g. DPT [31], GANs [10] and 2D diffusion models for image generation [7].
@@
To tackle such a challenge, one set of works require a time-consuming iterative optimization for obtaining novel view images though building a per-object model such as RealFusion [21], NeuralLift [49], SSDNerf [30], where the input image needs to be involved in a multi-round training process. We name them as training-based methods. Another set of works target at fast and efficient synthesis, which can finish the task with a single forward process of the given image, e.g. MCC [47], Graf [35], and Zero123 [18] , where we name then as training-free methods. In this paper, our method falls in the second category since more efficiency implies wider range of applications. In principle, the effectiveness and generalization of training-free methods comes from using a deep network to absorb the knowledge from a sufficient large amount of 2D/3D dataset. For example, one may adopt a large transformer [32, 47] or a diffusion-based models [2, 18] for direct synthesis.
@@
Among all of these training-free works, we are particularly interested in Zero123 [18], since it is the latest model yielding SoTA NVS quality, and generalized well to out-of-distribution images. We will discuss other models in related works in Sec. 2. Specifically, Zero123 finetunes a pre-trained image-to-image diffusion model [1] on current largest CAD dataset [6] with a novel embeded camera embedding inside the model. Although the quality are impressive, as shown in top-rows of Fig. 1, we found the results are still not geometrically consistent. While we think the issue could be largely mitigated or even solved if given enough training data such as 10x bigger than objaverse-XL [5], and adopted even larger pretrained models such as Stable-Diffusion-XL [28], the current model for NVS is still resource constrained, where the design of a model architecture is essential for producing better synthesized images.
@@
Therefore, we take a further step towards the architecture optimization by designing a model that handles the challenges from reconstruction and generation separately, which significantly improve the geometric consistency, as shown in the bottom-rows of Fig. 1. To motivate our architecture design, we did an in-depth study of the model in Zero123, and come to a conclusion that the inconsistency comes from the inherent uncertainty of the diffusion-based model where a non-deterministic process is performed given a noisy input. Therefore, it is difficult to maintain the 3D consistency across independently sampled views. To tackle this issue, we introduce a two stage model, as shown in Fig. 2. In the first stage, we setup a deterministic Scene Representation Transformer (SRT) which quickly transforms the observed image to novel views which roughly represents the object shape and appearances. In the second stage, we design a View-conditioned diffusion model, which is conditioned not only on camera embedding, but also the outputs from SRT, to produce clear and highly detailed hallucinated unseen images. Inside the models, to enforce 3D consistency, we additionally employ epipolor-guided attention and multi-view attention to better aggregate multi-view information. Finally, a hierarchy generation paradigm is Incorporated to generate long sequences of consistent 3D views, allowing a full experience of 360 observation of the provided object image. We name this architecture as Consistent 1-to-3, and experiment it over several popular datasets [6]. In all cases, our results outperform other SoTA methods in-term of both quality and consistency, demonstrating the effectiveness of the proposed method. Finally, Consistent 1-to-3 can take few-shot images as input, which could further improve the NVS performance.
@@
In summary, in this paper, we contribute a novel geometric-aware deep model, namely Consistent 1-to-3, for efficient novel view image synthesis. Based on experiments over multiple large datasets, it produces the state-of-the-art synthesis quality and geometry consistency across various camera view points. Our code and model will be released upon the acceptance of this paper.
@@ 2 Related Work
In this section, we briefly review the literature in novel-view synthesis, and summarize a method and attribute configuration comparison in Tab. 1. In the following, we will walk-through these works on NVS to show the uniqueness of Consistent 1-to-3. Notice we recognize the breadth and diversity within NVS by including some representative works in our discussion. While we acknowledge the existence of many similar and significant contributions, it is beyond our scope to cite them all.
@@
Single Instance Training-based Reconstruction As discussed in the introduction (Sec. 1), this thread of works introduce a multi-round optimization strategy to obtain an instance specific model for view synthesis. NeRF [23] first propose a neural representation for optimization with around 100 images, while produce realistic synthesised images. VolSDF [50] can produce nerf model with 10s of images by incorporating a 3D volume representation. Later works tries to further reduce the number of provided images, while maintaining the original renderring quality. RegNeRF [27] takes 3 images and adopt a geometry loss from depth patches. SinNerf [48], RealFusion [21], NeuralLift [49] further try to reduce the required images to only a single one by combining either depth map or Score Distillation Sampling [29] during the nerf model training. However, all these strategies needs a long-term training optimization, e.g. $\geq$ 30min for a single model for RealFusion, which hinders a wide range of interactive applications.
@@
Re-projection methods. This thread of works consider adopting an explicit geometry representation inside, and then do a 3D re-projection for novel view synthesis. IBRNet [44] and GPNR [40] adopts image-based view synthesis with a sparse set of nearby views with pixel or patch based embedding. PixelNeRF [51] trained a generalizable Nerf model with 13 object categories in ShapeNet while not showing its ability to large amount of instances in open-set images such as the ones in Objaverse [6]. NerFormer [32] further enlarges the 3D datasets to 50 categories with 18,619 instances, and proposes a transformer that can do view synthesis from a small number of images, later works like, MCC [47] introduce a volumn representation yielding better performance when a single image is given, while the rendered images are lacking details.
In addition to category-agnostic approaches, some works adopt categroy specific mesh models as priori to improve the results such as Human View Extrapolation (HumanVE) [55] or Pet Novel View Synthesis (PetNVS) [37]. These models can hardly generalize to open-set instances.
@@
Latent methods. This thread of work generalize the explicit pixel representation to latent ones, and learn MLP to directly decode the latent code to novel-view rather than performing the ray aggregation from NeRF, yielding much faster rendering speed, which is first proposed in LFN [39]. Later, SRT [34] further reduced the necessary observed images to few-shot. However, these methods have not show their generalization ability across open-set categories.
@@
Diffusion Models for NVS. This thread of work extends existing image diffusion models to conditional diffusion models where input views and target camera parameters are used as the condition for generating a target view. Watsonet al. [46] first proposed to use diffusion models for such a task by training a diffusion model on the SRN ShapeNet dataset [38]. SparseFusion [54] combines stable diffusion model [33] with epipolar feature transformer for a view-conditioned diffusion model in the latent space. GeNVS [2] proposed to improve the view consistency of diffusion model by re-projecting the latent feature of input view before diffusion denoising. However, all of these work are limited to their training data and have not been proved to generalize to arbitrary image input. Zero-1-to-3 [18] utilizes pre-trained image variation stable diffusion model [1] and fine-tunes it on a large 3D render dataset. It is found that such a fine-tuned model is able to generate high quality novel view output with relatively good generalizability compared to prior works.
@@
3D Diffusion Models. Utilizing categorical 3D datasets, numerous works have leveraged the diffusion model directly on 3D shapes. Zhou et al. [53] and Zeng et al. [43] employ a diffusion model for 3D point clouds. $PC^{2}$ [22] learns a point cloud generation model conditioned on a single image. Diffusion models based on various 3D representations, such as meshes [19], volumes [3], and neural implicit representations [11, 4], have also been explored. We have also seen rapid progress on the more challenging open-set (or universal) 3D generation [52] thanks to the release of Objaverse [6], a large-scale 3D model repository. Researchers from OpenAI also proposed Point-e [26], and Shap-e [13] for open-set 3D generation conditioned on texts or images using internal 3D datasets.
@@ 3 Method
Figure 2: Pipeline of Consistent-1-to-3. Given a single or a sparse set of input images, the encoder within the Scene Representation Transformer (SRT) translates the image(s) into a latent scene representation, effectively capturing implicit 3D information. The image rendering unfolds in two stages. The initial stage produces a rough yet geometry-gruonded output by cross-attending the queried pixels to the latent scene representation. Subsequently, these intermediate outputs are taken as input by the view-conditioned diffusion model, resulting in the visually appealing images that exhibit consistency with both the input images and the generated images from different viewpoints.
Given single observations of an object, denoted as $\boldsymbol{x}^{\mathrm{source}}$, with known camera parameters $\boldsymbol{\pi}^{\mathrm{source}}$, our goal is to synthesize a set of multi-view consistent images using target camera parameters $\boldsymbol{\pi}^{\mathrm{target}}$. As illustrated in Fig. 2, our pipeline consists of two key components: 1) A Scene Representation Transformer that learns a latent 3D representation given a single view; and 2) a view-conditioned diffusion model to capture the generative aspects of novel view synthesis. We first detail the pipeline in Sec. 3.1, and introduce several essential techniques to further enhance multi-view consistency in  Sec. 3.1. Lastly, we extend the proposed Novel View Synthesis pipeline to facilitate 3D reconstruction ( Sec. 3.2).

### v2

@@ Abstract
Zero-shot novel view synthesis (NVS) from a single image is an essential problem in 3D object understanding. While recent approaches that leverage pre-trained generative models can synthesize high-quality novel views from in-the-wild inputs, they still struggle to maintain 3D consistency across different views. In this paper, we present Consistent-1-to-3, which is a generative framework that significantly mitigates this issue. Specifically, we decompose the NVS task into two stages: (i) transforming observed regions to a novel view, and (ii) hallucinating unseen regions. We design a scene representation transformer and view-conditioned diffusion model for performing these two stages respectively. Inside the models, to enforce 3D consistency, we propose to employ epipolar-guided attention to incorporate geometry constraints, and multi-view attention to better aggregate multi-view information. Finally, we design a hierarchy generation paradigm to generate long sequences of consistent views, allowing a full $360^{\circ}$ observation of the provided object image. Qualitative and quantitative evaluation over multiple datasets demonstrates the effectiveness of the proposed mechanisms against state-of-the-art approaches. Our project page is at https://jianglongye.com/consistent123/.
@@ 1 Introduction
1) Single-view $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2) Sparse-views (2-4) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
3) Multi-view consistent $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
4) Generate unseen $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
5) Open-set generalization $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
6) Train-free for new instances $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 1: Comparison with prior methods. The rows indicate whether each method: 1) works with a single input view, 2) works with sparse (2-4) input views, 3) generate consistent multiple views, 4) hallucinates unseen regions, 5) generalizes to open-set instances instead of limited categories, 6) free of time-consuming training for new instances. Our proposed method stands out as the only approach that possesses all these advantages.
@@
Synthesizing high-quality and visually consistent novel views of real-world objects from a single input image has been a long-standing problem in computer vision. Its applications are widespread, ranging from content creation [20, 33], robotic manipulation and navigation [29], to AR/VR [16].
@@
This task is challenging as it involves difficulties from both reconstruction and generation, which require not only accurate geometry transformation of observed parts but also hallucination of unseen regions. In recent years, the field has seen significant progress with the advance of large-scale data-driven deep 2D models. For example, transformer models for single view reconstruction, e.g., DPT [35], GANs [11] and 2D diffusion models for image generation [8].
@@
To tackle such a challenge, one set of works requires a time-consuming iterative optimization to obtain novel view images through building an object-specific model such as RealFusion [25], NeuralLift [55] and SSDNerf [34], where the input image needs to be involved in a multi-round training process. We note them as training-based methods. Another set of works targets fast and efficient synthesis, which can complete the task with a single forward process of the given image, e.g., MCC [53], Graf [40] and Zero123 [22]. We note them as training-free methods. In this paper, our method falls into the second category since more efficiency implies a wider range of applications. In principle, the effectiveness and generalization of training-free methods come from using a deep network to learn the prior information from a sufficiently large amount of 2D/3D data. For example, one may adopt large transformers [36, 53] or diffusion-based models [2, 22] for view synthesis.
@@
Among training-free methods, we are particularly interested in Zero123 [22], since it is among the recent models that can produce SoTA NVS quality, and generalize well to out-of-distribution images. We will discuss other models in related works in Sec. 2. Specifically, Zero123 finetunes a pre-trained image-to-image diffusion model [1] on a large CAD dataset [7] with a novel camera embedding condition in the diffusion model. Although the quality is impressive, as shown in the top rows of Fig. 1, we found the results are still not geometrically consistent. While we conjecture this issue could be largely mitigated if given enough training data such as 10x bigger than objaverse-XL [6] and adopted even larger pre-trained models such as Stable-Diffusion-XL [32], the current model for NVS is still resource-constrained, and the architecture design is essential for the quality of synthesized images.
@@
Therefore, we take a further step towards architecture optimization by designing a model that handles the challenges from reconstruction and generation separately, which significantly improves the geometric consistency, as shown in the bottom rows of Fig. 1. To motivate our architecture design, we did an in-depth study of the Zero123 model, and concluded that the inconsistency comes from the inherent uncertainty of the diffusion-based model where a non-deterministic process is performed given a noisy input. Therefore, it is difficult to maintain the 3D consistency across independently sampled views. To tackle this issue, we introduce a two-stage model, as shown in Fig. 2. In the first stage, we set up a deterministic Scene Representation Transformer (SRT) [39] which quickly transforms the observed image into novel views that roughly represent the object’s shape and appearance. In the second stage, we design a view-conditioned diffusion model, which is conditioned not only on the camera embedding but also the outputs from SRT, to produce clear and detailed images for unseen regions. To enforce 3D consistency, we additionally employ epipolar-guided attention [48] and multi-view attention [42] to better aggregate multi-view information. Finally, a hierarchy generation paradigm is proposed to generate long sequences of consistent 3D views, allowing a full experience of $360^{\circ}$ observation of the provided object image. We name this architecture as Consistent 1-to-3 and experiment with it over several popular datasets [7]. In all cases, our results outperform other SoTA methods in terms of both quality and consistency, demonstrating the effectiveness of the proposed method. Finally, Consistent 1-to-3 can further improve the NVS performance when using few-shot images as input.
@@
In summary, our paper contributes a novel geometric-aware deep model, namely Consistent 1-to-3, for efficient novel view synthesis. With extensive experiments on multiple datasets, it produces state-of-the-art synthesis quality and geometric consistency across various camera viewpoints.
@@ 2 Related Work
In this section, we briefly review the literature in novel-view synthesis and group related methods based on their attributes, shown in Tab. 1. In the following, we will walk through these works on NVS to show the uniqueness of Consistent 1-to-3. We recognize the breadth and diversity within NVS, and only include representative works due to limited space. While we acknowledge the existence of other significant contributions, citing them all is beyond our scope.
@@
Single Instance Training-based Reconstruction. As discussed in Sec. 1, this thread of works introduces a multi-round optimization strategy to obtain an instance-specific model for view synthesis. NeRF [27] first proposes a neural network that optimizes around 100 images to produce realistic synthesized images. VolSDF [56] can train a NeRF model with tens of images by incorporating a 3D volume representation. Later works try to further reduce the number of required images, while maintaining the original rendering quality. RegNeRF [31] takes 3 images and adopts a geometry loss from depth patches. SinNerf [54], RealFusion [25] and NeuralLift [55] further reduce the required images to a single one by combining either depth map or Score Distillation Sampling [33] during the nerf model training. However, all these strategies need time-consuming optimization, e.g., more than 30 mins for a single model for RealFusion, which hinders a wide range of interactive applications.
@@
Re-projection methods. This thread of works considers adopting an explicit geometry representation, then doing a 3D re-projection for novel view synthesis. IBRNet [50] and GPNR [46] adopts image-based view synthesis using a sparse set of nearby views with pixel or patch-based embedding. PixelNeRF [57] trained a generalizable Nerf model with 13 object categories from ShapeNet [3] while not showing its ability to learn from large amount of instances in open-set images such as Objaverse [7]. NerFormer [36] further enlarges the 3D datasets to 50 categories with 18,619 instances and proposes a transformer for view synthesis from a small number of images. Later, MCC [53] introduce a volume representation yielding better performance when a single image is given, but the rendered images lack details. In addition to category-agnostic approaches, some works adopt category-specific mesh models as prior to improve the results such as Human View Extrapolation (HumanVE) [61] or Pet Novel View Synthesis (PetNVS) [43]. These models can hardly generalize to open-set instances.
@@
Latent methods. This thread of work generalizes the explicit pixel representation to latent ones and learns MLP to directly decode the latent code to novel-view rather than performing ray aggregation from NeRF, yielding much faster rendering speed. This is first proposed in LFN [45]. Later, SRT [39] further reduced the necessary observed images to a few shots. However, these methods have not shown their generalization ability across open-set categories.
@@
Diffusion Models for NVS. This thread of work extends existing image diffusion models to conditional diffusion models where input views and target camera parameters are used as the conditions for generating a target view. Watson et al. [52] first proposed to use diffusion models for such a task by training a diffusion model on the SRN ShapeNet dataset [44]. SparseFusion [60] combines stable diffusion model [37] with epipolar feature transformer for a view-conditioned diffusion model in the latent space. GeNVS [2] proposed to improve the view consistency of the diffusion model by re-projecting the latent feature of input view before diffusion denoising. However, all of these works are limited to their training data and have not been proven to generalize to arbitrary image input. Zero-1-to-3 [22] utilizes pre-trained stable diffusion model [1] and fine-tunes it on a large 3D render dataset. It is found that such a fine-tuned model is able to generate high-quality novel view output with relatively good generalizability compared to prior works. The follow-up work One-2-3-45 [21] takes advantage of the generalization capabilities of Zero-1-to-3 and employs a feed-forward network to obtain a mesh representation. More recently, MVDream [42] proposes a similar multi-view diffusion mechanism to generate geometrically consistent multi-view images from a given text prompt.
@@
3D Diffusion Models. Utilizing categorical 3D datasets, numerous works have leveraged diffusion models directly on 3D shapes. Zhou et al. [59] and Zeng et al. [49] employ diffusion models for 3D point clouds. $PC^{2}$ [26] learns a point cloud generation model conditioned on a single image. Diffusion models based on various 3D representations, such as meshes [23], volumes [4], and neural implicit representations [13, 5], have also been explored. We have also seen rapid progress on the more challenging open-set (or universal) 3D generation [58] thanks to the release of Objaverse [7], a large-scale 3D model repository. Researchers from OpenAI also proposed Point-e [30], and Shap-e [15] for open-set 3D generation conditioned on texts or images using internal 3D datasets.
@@ 3 Method
Figure 2: Pipeline of Consistent-1-to-3. Given a single or a sparse set of input images, the encoder within Scene Representation Transformer (SRT) translates the image(s) into latent scene representations, effectively capturing implicit 3D information. The image rendering unfolds in two stages. The initial stage produces a rough yet geometry-grounded output by cross-attending the queried pixels to the latent scene representation. Subsequently, these intermediate outputs are taken as input by the view-conditioned diffusion model, resulting in visually appealing images that exhibit consistency among the input and generated images from different viewpoints.
Given a single observation of an object, denoted as $\boldsymbol{x}^{\mathrm{source}}$, with known camera parameters $\boldsymbol{\pi}^{\mathrm{source}}$, our goal is to synthesize a set of multi-view consistent images using target camera parameters $\boldsymbol{\pi}^{\mathrm{target}}$. As illustrated in Fig. 2, our pipeline consists of two key components: 1) A Scene Representation Transformer that learns a latent 3D representation given a single view; and 2) a view-conditioned diffusion model to capture the generative aspects of novel view synthesis. We first detail the pipeline in Sec. 3.1, and introduce several essential techniques to further enhance multi-view consistency in  Sec. 3.2. Lastly, we extend the proposed Novel View Synthesis pipeline to facilitate 3D reconstruction ( Sec. 3.3).